# A Nonabelian Hodge Correspondence for Principal Bundles in Positive Characteristic

### v1

@@ Abstract.
In this paper, we prove a nonabelian Hodge correspondence for principal bundles on a smooth variety $X$ in positive characteristic, which generalizes the Ogus-Vologodsky correspondence for vector bundles. Then we extend the correspondence to logahoric torsors over a log pair $(X,D)$, where $D$ a reduced normal crossing divisor in $X$. As an intermediate step, we prove a correspondence between principal bundles on root stacks $\mathscr{X}$ and parahoric torsors on $(X,D)$, which generalizes the correspondence on curves given by Balaji–Seshadri to higher dimensional case.
@@ 1. Introduction
Ogus–Vologodsky established the nonabelian Hodge correspondence for $G={\rm GL}_{n}$ in positive characteristic [OV07]. A natural question is how to generalize Ogus–Vologodsky’s result to an arbitrary reductive group $G$. On the other hand, Chen–Zhu established a nonabelian Hodge correspondence (not in one to one manner) for reductive groups on curves with the help of Hitchin morphisms [CZ15]. For nilpotent objects with small exponent their correspondence coincides with the Ogus-Vologodsky correspondence. However, their approach seems to be hard to be generalized to higher dimensional varieties, because there is still a lack of numerous necessary properties of the Hitchin morphism (see [CN20, Conjecture 3.2 and Conjecture 5.2]).
@@
As remarked in [LSZ15, Remark 2.4], the exponential twisting approach to the Ogus-Vologodsky correspondence should respect $G$-structure for any subgroup $G\subset\mathrm{GL}$. In this paper, we aim to work out this point in detail, which is going to be applied to construct many absolutely irreducible crystalline $G$-local systems in a subsequent work. As would be clear to the reader, the key point in this approach is the construction of exponential maps for reductive groups with good properties. Fortunately, this has been studied by Serre [Ser94] and several other authors in a series work [Sei00, McN02, Sob15]. In particular, Sobaje gave a definition of the exponential map in positive characteristic [Sob15, Definition 2.3] and prove its existence and uniqueness when the characteristic is pretty good [Sob15, Theorem 3.4]. These knowledge suffice for generalizing the result of Lan–Sheng–Zuo [LSZ15] in the vector bundle case to principal $G$-bundles, where $G$ denotes for a connected split reductive group $G$ over a perfect field $k$ in char $p>0$. In effect, we establish a correspondence between nilpotent $G$-Higgs bundles and nilpotent flat $G$-bundles of exponent $\leq p-1$ for a smooth variety over $k$ which is $W_{2}(k)$-liftable. Next, we generalize this correspondence to the special log setting $(X,D)$, where $X$ is a smooth variety and $D$ is a reduced normal crossing divisor of $X$. There has been the work by Schepler [Sch05] and the work [LSYZ19] via exponential twisting in the vector bundle case, and the work by Krishnamoorthy and the first named author [KS20] for parabolic bundles over curves. The right generalization to principal bundles of parabolic Higgs bundles which have simple poles along $D$ is the notion of logahoric (=logarithmic parahoric) Higgs torsors, which was proposed by P. Boalch [Boa11] in characteristic zero (see also [KSZ24]). Like in [KS20], we first establish a logarithmic version of the correspondence over a tame root stack. However, the procedure of descending logarithmic Higgs torsors/connections from root stacks to logahoric Higgs torsors/connections over $(X,D)$ is bit different from the generalized Biswas-Iyer-Simpson correspondence for parabolic vector bundles in loc. cit. When $X$ is projective equipped with an ample line bundle $H$, one has the notion of R-stability [Ram75, Ram96a] for principal bundles which generalizes the Mumford-Takemoto slope stability in vector bundle case [Ram75, §3]. It has been further generalized to a log pair $(X,D)$ in [KSZ24]. We show that our correspondence preserves the R-stability condition, as one may expect.
@@
Now we give the setup and main results in this paper. Let $k$ be a perfect field in positive characteristic $p$. Let $X$ be a $W_{2}(k)$-liftable smooth variety over $k$ and denote by $X^{\prime}$ the Frobenius pullback. Let $G$ be a connected split reductive linear algebraic group over $k$ and we suppose that the characteristic $p$ is pretty good [Her13, Definition 2.11]. We introduce the following categories
  • ${\rm HIG}_{p-1}(X^{\prime}/k,G)$: the category of nilpotent $G$-Higgs bundles of exponent $\leq p-1$ on $X$;
@@
and prove that they are equivalent by establishing the exponential twisting of the Cartier and inverse Cartier transform in §3.
@@
Suppose $X$ is projective and we fix an ample line bundle $H$ on $X$. We follow Ramanathan’s approach [Ram75, Ram96a] to define the stability condition of $G$-bundles on $X$ with an arbitrary dimension, which is called $R$-stability condition (see §2.3). Denote by ${\rm HIG}^{s(s)}_{p-1}(X^{\prime}/k,G)\subseteq{\rm HIG}_{p-1}(X^{\prime}/k)$ (resp. ${\rm MIC}^{s(s)}_{p-1}(X/k,G)\subseteq{\rm MIC}_{p-1}(X/k)$) the full subcategory of $R$-(semi)stable nilpotent $G$-Higgs bundles (resp. flat $G$-bundles) of exponent $\leq p-1$. We prove that the correspondence also holds under $R$-stability conditions.
Theorem 1.2 (Theorem 3.6).
The categories ${\rm HIG}^{s(s)}_{p-1}(X^{\prime}/k,G)$ and ${\rm MIC}^{s(s)}_{p-1}(X/k,G)$ are equivalent.
Now we consider the correspondence on $X$ together with a reduced normal crossing divisor $D=\sum_{i=1}^{s}D_{i}$, where $D_{i}$ are irreducible components for $1\leq i\leq s$. Following Simpson’s idea [Sim90], the analogue of principal bundles on $X\backslash D$ is parahoric torsors [HKSZ22], especially when discussing stability conditions. In a nutshell, we prove the equivalence of categories between nilpotent logahoric Higgs torsors and nilpotent logahoric connections of exponent $\leq p-1$. Here is a sketch of the proof. We equip each $D_{i}$ with a tame weight $\alpha_{i}$, where a tame weight is a cocharacter with rational coefficients such that the denominator $d_{i}$ is coprime to $p$. Denote by $\boldsymbol{\alpha}$ (resp. $\boldsymbol{d}$) the collection of weights (resp. integers). With respect to the above data, people constructed a (tame) root stack $\mathscr{X}$ ([MO05, Theorem 4.1]) and a smooth affine group scheme $\mathcal{G}_{\boldsymbol{\alpha}}$ on $X$ [BP22, Theorem 5.4]. We define the following categories
  • ${\rm HIG}_{p-1}(\mathscr{X}_{\rm log}/k,G)$: the category of nilpotent logarithmic $G$-Higgs bundles of exponent $\leq p-1$ on $\mathscr{X}$;
@@
Similar to the case of curves [BS15, §3], we prove that parahoric $\mathcal{G}_{\boldsymbol{\alpha}}$-torsors on $X$ are in one-to-one correpsondence with $G$-bundles on $\mathscr{X}$ of type $\boldsymbol{\rho}$, where $\boldsymbol{\rho}=\{\rho_{i},1\leq i\leq s\}$ is a collection of representations $\rho_{i}:\mu_{d_{i}}\rightarrow T$ determined by $\boldsymbol{\alpha}$ and $T$ is a maximal torus of $G$.
@@
This correspondence carries over $G$-bundles with Higgs field/connection. It has been done in [KSZ24] for curves in characteristic zero, and in [KS20] for vector bundles over curves in positive characteristic. We first introduce some definitions. A logahoric $\mathcal{G}_{\boldsymbol{\alpha}}$-Higgs torsor is a pair $(\mathcal{E},\vartheta)$, where $\mathcal{E}$ is a parahoric $\mathcal{G}_{\boldsymbol{\alpha}}$-torsor and $\vartheta:X\rightarrow\mathcal{E}(\mathfrak{g})\otimes\Omega_{X}({\rm log}\,D)$ is a logarithmic Higgs field. A logahoric $\mathcal{G}_{\boldsymbol{\alpha}}$-connection is a pair $(\mathcal{V},\nabla^{\prime})$, where $\mathcal{V}$ is a parahoric $\mathcal{G}_{\boldsymbol{\alpha}}$-torsor and $\nabla^{\prime}:\mathcal{O}_{\mathcal{V}}\rightarrow\mathcal{O}_{\mathcal{V}}\otimes\Omega_{X}({\rm log}\,D)$ is a logarithmic integrable connection. We define
@@
Note that, in the above definition, the residue of a logarithmic connection on a parahoric torsor is not nilpotent and the semisimple part is given by the weights $\boldsymbol{\alpha}$.
@@
Furthermore, the above equivalences also hold under $R$-stability conditions for Higgs bundles and flat $G$-bundles. Therefore, we follow the diagram below and prove a logahoric version of nonabelian Hodge correspondence:
@@
Suppose that $(X,D)$ is $W_{2}(k)$-liftable. The categories ${\rm HIG}_{p-1}(X^{\prime}_{\rm log}/k,\mathcal{G}_{\boldsymbol{\alpha}})$ and ${\rm MIC}_{p-1}(X_{\rm log}/k,\mathcal{G}_{\boldsymbol{\alpha}^{\prime}})$. Moreover, the equivalence preserves stability conditions.
@@
Ogus–Vologodsky established the nonabelian Hodge correspondence in positive characteristic for $G={\rm GL}_{n}$ [OV07] and the logarithmic case was studied by Schepler [Sch05]. Our work is a generalization of Ogus–Vologodsky’s correspondence for general reductive groups and we also extend the results to parahoric torsors. Also, Chen–Zhu introduced an analogue of Hitchin’s equations for principal bundles in positive characteristic
$\begin{cases}\psi(\nabla-\theta)=0\\ ((\nabla-\theta)\otimes\nabla^{can})(\psi(\nabla))=0\end{cases}.$
@@
Parahoric torsors were studied by many groups of people. Heinloth studied uniformization property of parahoric torsors [Hei10]. After that, Balaji–Seshadri constructed the moduli space of parahorc torsors in the case of curves [BS15]. Recently, Balaji–Pandey defined parahoric group schemes over higher dimensional bases [BP22]. Similar to the case of curves [BS15, §3], parahoric torsors are equivalent to principal bundles on an appropriate root stacks. This property is observed by Balaji–Pandey locally [BP22] and is proven in §5.2.
The $R$-stability condition for $G$-bundles is introduced by Ramanathan in the case of curves [Ram75, Ram96a]. When $G$ is a classical group, the $R$-stability condition of a principal bundle is equivalent to the slope stability condition of the corresponding associated bundle in the case of curves. We refer the reader to [CS21, §6] and [Yan22] for more details. In a similar way, Heinloth define a stability condition for parahoric torsors [Hei17, §1]. The stability condition he considered is equivalent to the slope stability condition, which does not include the information of weights. On the other hand, the stability condition considered for parahoric torsors in this paper include the information of weights (Definition 5.6 and 5.7), which is equivalent to the stability condition of the corresponding parabolic bundles when $G={\rm GL}_{n}$ [KSZ24, §5]. Moreover, it is also equivalent to the one given by Balaji-Seshadri [BS15]. Moreover, as proved in this paper, the stability condition of parahoric torsors is equivalent to the stability condition of the corresponding $G$-bundles on root stacks (Proposition 5.9).
In the parahoric case, the residues of nilpotent logarithmic connections are not nilpotent in general and the semisimple part of the residue is given by the corresponding weight. This phenomenon actually comes from the establishment of the equivalence between parahoric torsors on $X$ and principal bundles on the root stack $\mathscr{X}$. We introduce a new element $\Delta_{i}$ for each $1\leq i\leq s$ and define new transition functions. The change of transition functions results in the change of residues. In positive characteristic, this phenomenon first appears in [KS20], where the authors studied the case of $G={\rm GL}_{n}$ and named the condition as adjustedness [KS20, Definition 2.9]. Later on, the authors generalized the result to parahoric torsors on curves [LS24], which also generalized Chen–Zhu’s result [CZ15]. Moreover, in characteristic zero, a similar discussion is also given in [Sim90, §5].
@@ 2. Preliminaries
In this section, we introduce the main objects considered in this paper: (semistable) nilpotent $G$-Higgs bundles and (semistable) nilpotent flat $G$-bundles. In §2.1, we review definitions of $G$-Higgs bundles and flat $G$-bundles, and a very good reference is [CZ15, Appendix]. In §2.2, we introduce the nilpotency and review the exponential map for linear algebraic groups given in [Sob15]. In §2.3, we give the definition of stability conditions, which come from Ramanathan’s definition for principal bundles [Ram75, Ram96a] and are called $R$-stability conditions. We also refer the reader to [KSZ24] for more details about the case in characteristic zero.
@@ 2.1. G-Higgs bundles and Flat G-bundles
Let $X$ be a smooth variety over $k$, which is a perfect field in positive characteristic $p$. Denote by $T_{X/k}$ (resp. $\Omega_{X/k}$) the relative tangent (resp. cotangent) sheaf. Sometimes we use the notation $T_{X}$ and $\Omega_{X}$ if there is no ambiguity. Since we are working in characteristic $p$, we have the following well-known diagram
@@
  • $X^{\prime}:=X\times_{{\rm Spec}\,k,F}{\rm Spec}\,k$,
@@
Let $G$ be a connected split reductive linear algebraic group over $k$ with maximal torus $T$. Denote by $\mathcal{R}$ the set of roots. Let $\mathfrak{g}$ (resp. $\mathfrak{t}$) be the Lie algebra of $G$ (resp. $T$). Let $E$ be a $G$-bundle on $X$ with adjoint bundle $E(\mathfrak{g}):=E\times_{G}g$. The Lie bracket [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] of $\mathfrak{g}$ induces a well-defined binary operation
$E(\mathfrak{g})\otimes_{\mathcal{O}_{X}}E(\mathfrak{g})\rightarrow E(\mathfrak{g}),$
and we use the same notation [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] for this morphism (see [BBP17, §8] for instance). Furthermore, if the tensor product is defined over $\mathcal{O}_{X}$, we use the notation $E(\mathfrak{g})\otimes E(\mathfrak{g})$ for convenience. Now given a section $\theta\in H^{0}(X,E(\mathfrak{g})\otimes\Omega_{X})$, the composition of the following maps
$X\xrightarrow{\theta\otimes\theta}(E(\mathfrak{g})\otimes\Omega_{X})\otimes(E(\mathfrak{g})\otimes\Omega_{X})\rightarrow(E(\mathfrak{g})\otimes E(\mathfrak{g})\otimes(\Omega_{X}\otimes\Omega_{X})\xrightarrow{[\cdot,\cdot]\otimes\wedge}\rightarrow E(\mathfrak{g})\otimes\Omega^{2}_{X}$
is denoted by $\theta\wedge\theta$. We say that $\theta$ is integrable if $\theta\wedge\theta=0$, and an integrable section is called a Higgs field.
@@
A $G$-Higgs bundle on $X$ is a pair $(E,\theta)$, where $E$ is a $G$-bundle on $X$ and $\theta$ is a Higgs field.
Now we consider connections on $G$-bundles. There is a canonical connection on $\mathcal{O}_{G\times X}$ and denote it by $\nabla_{G}$. Let $V$ be a $G$-bundle. A (integrable) $G$-connection $\nabla$ on $V$ is a (integrable) connection $\nabla:\mathcal{O}_{V}\rightarrow\mathcal{O}_{V}\otimes_{\mathcal{O}_{X}}\Omega_{X}$ such that the following diagram commutes

### v2

@@ Abstract.
In this paper, we prove a nonabelian Hodge correspondence for principal bundles on a smooth variety $X$ in positive characteristic, which generalizes the Ogus–Vologodsky correspondence for vector bundles. Then we extend the correspondence to parahoric torsors over a log pair $(X,D)$, where $D$ a reduced normal crossing divisor in $X$. As an intermediate step, we prove a correspondence between principal bundles on root stacks $\mathscr{X}$ and parahoric torsors on $(X,D)$, which generalizes the correspondence on curves given by Balaji–Seshadri to higher dimensional case.
@@ 1. Introduction
Ogus–Vologodsky established the nonabelian Hodge correspondence for $G={\rm GL}_{n}$ in positive characteristic [OV07]. A natural question is how to generalize Ogus–Vologodsky’s result to an arbitrary reductive group $G$. Chen–Zhu established a nonabelian Hodge correspondence for reductive groups on curves with the help of Hitchin morphisms [CZ15]. For nilpotent objects with small exponent their correspondence coincides with the Ogus–Vologodsky correspondence. However, their approach seems to be hard to be generalized to higher dimensional varieties, because there is still a lack of numerous necessary properties of the Hitchin morphism (see [CN20, Conjecture 3.2 and Conjecture 5.2] for instance).
@@
As remarked in [LSZ15, Remark 2.4], the exponential twisting approach to the Ogus–Vologodsky correspondence should respect $G$-structure for any subgroup $G\subset\mathrm{GL}_{n}$. In this paper, we aim to work out this point in detail, which is going to be applied to construct many absolutely irreducible crystalline $G$-local systems in a subsequent work. As would be clear to the reader, the key point in this approach is the construction of exponential maps for reductive groups with good properties. Fortunately, this has been studied by Serre [Ser94] and several other authors in a series work [Sei00, McN02, Sob15]. In particular, Sobaje proved its existence and uniqueness when the characteristic is good [Sob15, Theorem 3.4]. These knowledge suffice for generalizing the result of Lan–Sheng–Zuo [LSZ15] in the vector bundle case to principal $G$-bundles, where $G$ denotes for a connected split reductive group $G$ over a perfect field $k$ in positive characteristic $p$. In effect, we establish a correspondence between nilpotent $G$-Higgs bundles and nilpotent flat $G$-bundles of exponent $\leq p-1$ for a smooth variety over $k$ which is $W_{2}(k)$-liftable. Next, we generalize this correspondence to the special log setting $(X,D)$, where $X$ is a smooth variety and $D$ is a reduced normal crossing divisor of $X$. There has been the work by Schepler [Sch05] and the work [LSYZ19] via exponential twisting in the vector bundle case, and the work by Krishnamoorthy and the first named author [KS20] for parabolic bundles over curves. The right generalization to principal bundles of parabolic Higgs bundles which have simple poles along $D$ is the notion of logahoric (=logarithmic parahoric) Higgs torsors (see [KSZ24] for instance). Like in [KS20], we first establish a logarithmic version of the correspondence over a tame root stack. However, the procedure of descending logarithmic Higgs torsors/connections from root stacks to logahoric Higgs torsors/connections over $(X,D)$ is a bit different from the generalized Biswas-Iyer-Simpson correspondence for parabolic vector bundles in loc. cit. When $X$ is projective equipped with an ample line bundle $H$, one has the notion of $R$-stability [Ram75, Ram79] for principal bundles which generalizes the Mumford-Takemoto slope stability in vector bundle case [Ram75, §3]. It has been further generalized to a log pair $(X,D)$ in [KSZ24]. We show that our correspondence preserves the $R$-stability condition, as one may expect.
@@
Now we give the setup and main results in this paper. Let $k$ be a perfect field in positive characteristic $p$. Let $X$ be a $W_{2}(k)$-liftable smooth variety over $k$ and denote by $X^{\prime}$ the Frobenius pullback. Let $G$ be a connected split reductive linear algebraic group over $k$ and we suppose that the characteristic $p$ is good [Her13, §2.9]. We introduce the following categories
  • ${\rm HIG}_{p-1}(X^{\prime}/k,G)$: the category of nilpotent $G$-Higgs bundles of exponent $\leq p-1$ on $X^{\prime}$;
@@
and prove that they are equivalent by establishing the Cartier and inverse Cartier transform via exponential twisting in §3.
@@
Suppose $X$ is projective and we fix an ample line bundle $H$ on $X$. We follow Ramanathan’s approach [Ram75, Ram96a] to define the stability condition of $G$-bundles on $X$ of an arbitrary dimension, which is called $R$-stability condition (see §2.1). Denote by ${\rm HIG}^{s(s)}_{p-1}(X^{\prime}/k,G)\subseteq{\rm HIG}_{p-1}(X^{\prime}/k)$ (resp. ${\rm MIC}^{s(s)}_{p-1}(X/k,G)\subseteq{\rm MIC}_{p-1}(X/k)$) the full subcategory of $R$-(semi)stable nilpotent $G$-Higgs bundles (resp. flat $G$-bundles) of exponent $\leq p-1$. We prove that the correspondence also holds under $R$-stability conditions (Theorem 3.8).
Now we consider the correspondence on $X$ together with a reduced normal crossing divisor $D=\sum_{i=1}^{s}D_{i}$, where $D_{i}$ are irreducible components for $1\leq i\leq s$. By introducing parahoric torsors, we prove a logarithmic version of the nonabelian Hodge correspondence. In a nutshell, we prove the equivalence of categories between nilpotent logahoric Higgs torsors and nilpotent logahoric connections of exponent $\leq p-1$. Here is a sketch of the proof. We equip each $D_{i}$ with a tame weight $\alpha_{i}$, where a tame weight is a cocharacter with rational coefficients such that the denominator $d_{i}$ is coprime to $p$. Denote by $\boldsymbol{\alpha}$ (resp. $\boldsymbol{d}$) the collection of weights (resp. integers). With respect to the above data, there exists a (tame) root stack $\mathscr{X}$ (see [MO05, Theorem 4.1] for example). We define the following categories
  • ${\rm HIG}_{p-1}(\mathscr{X}^{\prime}_{\rm log}/k,G)$: the category of nilpotent logarithmic $G$-Higgs bundles of exponent $\leq p-1$ on $\mathscr{X}^{\prime}$;
@@
Then let $\mathscr{F}$ be a $G$-bundle on $\mathscr{X}$. We can construct a parahoric group scheme $\mathcal{G}_{\boldsymbol{\alpha}}$ by Lemma 5.2. Note that this parahoric group scheme $\mathcal{G}_{\boldsymbol{\alpha}}$ depends on the given $G$-bundle $\mathscr{F}$ and the weights $\boldsymbol{\alpha}$ is determined by the type of $\mathscr{F}$. Similar to the case of curves [BS15, §3], we prove that parahoric $\mathcal{G}_{\boldsymbol{\alpha}}$-torsors on $X$ are in one-to-one correpsondence with $G$-bundles on $\mathscr{X}$ of type $\boldsymbol{\rho}$, where $\boldsymbol{\rho}=\{\rho_{i},1\leq i\leq s\}$ is a collection of representations $\rho_{i}:\mu_{d_{i}}\rightarrow T$ determined by $\boldsymbol{\alpha}$ and $T$ is a maximal torus of $G$.
@@
This correspondence carries over $G$-bundles with Higgs field/connection. It has been done in [KSZ24] for curves in characteristic zero, and in [KS20] for vector bundles over curves in positive characteristic. We first introduce some definitions. A logahoric $\mathcal{G}_{\boldsymbol{\alpha}}$-Higgs torsor is a pair $(\mathcal{E},\vartheta)$, where $\mathcal{E}$ is a parahoric $\mathcal{G}_{\boldsymbol{\alpha}}$-torsor and $\vartheta:X\rightarrow\mathcal{E}(\mathfrak{g})\otimes\Omega_{X}({\rm log}\,D)$ is a logarithmic Higgs field. A logahoric $\mathcal{G}_{\boldsymbol{\alpha}}$-connection is a pair $(\mathcal{V},\nabla^{\prime})$, where $\mathcal{V}$ is a parahoric $\mathcal{G}_{\boldsymbol{\alpha}}$-torsor and $\nabla^{\prime}:\mathcal{O}_{\mathcal{V}}\rightarrow\mathcal{O}_{\mathcal{V}}\otimes\Omega_{X}({\rm log}\,D)$ is a logarithmic integrable $\mathcal{G}_{\boldsymbol{\alpha}}$-connection. We define
@@
Note that, in the above definition, the residue of a logarithmic connection on a parahoric torsor is not necessarily to be nilpotent and the semisimple part is given by the weights $\boldsymbol{\alpha}$.
@@
Furthermore, the above equivalences also hold under $R$-stability conditions for $G$-Higgs bundles and flat $G$-bundles. Therefore, we follow the diagram below and prove a logahoric version of the nonabelian Hodge correspondence:
@@
Suppose that $(X,D)$ is $W_{2}(k)$-liftable. The categories ${\rm HIG}_{p-1}(X^{\prime}_{\rm log}/k,\mathcal{G}_{\boldsymbol{\alpha}})$ and ${\rm MIC}_{p-1}(X_{\rm log}/k,\mathcal{G}_{\boldsymbol{\alpha}^{\prime}})$ are equivalent. Moreover, the equivalence preserves stability conditions.
@@
Ogus–Vologodsky established the nonabelian Hodge correspondence in positive characteristic for $G={\rm GL}_{n}$ [OV07] and the logarithmic case was studied by Schepler [Sch05]. Our work is a generalization of Ogus–Vologodsky’s correspondence for reductive groups and we also extend the results to parahoric torsors. Also, Chen–Zhu introduced an analogue of Hitchin’s equations for principal bundles in positive characteristic
$\begin{cases}\psi(\nabla-\theta)=0\\ ((\nabla-\theta)\otimes\nabla^{can})(\psi(\nabla))=0\ ,\end{cases}$
@@
Parahoric torsors on curves were studied by many groups of people. Heinloth studied uniformization property of parahoric torsors [Hei10] and Yun studied parahoric Higgs torsors in [Yun11]. After that, Balaji–Seshadri constructed the moduli space of parahoric torsors in the case of curves [BS15]. For higher dimensional varieties, we give the definition of parahoric group schemes (Definition 5.1), which is inspired by Balaji–Pandey’s result [BP22, Theorem 5.4], and prove that parahoric torsors are equivalent to principal bundles on an appropriate root stack (Proposition 5.5).
The $R$-stability condition for $G$-bundles is introduced by Ramanathan in the case of curves [Ram75, Ram96a], and studied further in higher dimensional case [AB01, GLSS08]. When $G$ is a classical group, the $R$-stability condition of a principal bundle is equivalent to the slope stability condition of the corresponding associated bundle in the case of curves. We refer the reader to [CS21, §6] and [Yan22] for more details. In a similar way, Heinloth define a stability condition for parahoric torsors [Hei17, §1]. The stability condition he considered is equivalent to the slope stability condition, which does not include the information of weights. On the other hand, the stability condition considered for parahoric torsors in this paper is an analogue of that given by Balaji–Seshadri [BS15, §6] and includes the information of weights (Definition 5.9), which is equivalent to the stability condition of the corresponding parabolic bundles when $G={\rm GL}_{n}$.
In the logarithmic case, the residues of nilpotent logarithmic connections are not nilpotent in general and the semisimple part of the residue is given by the corresponding weight. This phenomenon actually comes from the establishment of the equivalence between parahoric torsors on $X$ and principal bundles on the root stack $\mathscr{X}$. In positive characteristic, this phenomenon first appears in [KS20], where the authors studied the case of $G={\rm GL}_{n}$ and named the condition as adjustedness [KS20, Definition 2.9]. Later on, the authors generalized the result to parahoric torsors on curves [LS24], which also generalized Chen–Zhu’s result [CZ15]. Moreover, in characteristic zero, a similar discussion is also given in [Sim90, §5].
@@ 2.1. G-Higgs bundles and flat G-bundles
Let $X$ be a smooth variety over a perfect field $k$ in positive characteristic $p$. Denote by $T_{X/k}$ (resp. $\Omega_{X/k}$) the relative tangent (resp. cotangent) sheaf. If there is no ambiguity, we use the notation $T_{X}$ and $\Omega_{X}$ for simplicity. We have the following well-known diagram about the Frobenius morphism:
@@
  • $X^{\prime}:=X\times_{{\rm Spec}\,k,F_{k}}{\rm Spec}\,k$,
@@
Now let $G$ be a connected split reductive linear algebraic group over $k$ with maximal torus $T$. Let $\mathfrak{g}$ (resp. $\mathfrak{t}$) be the Lie algebra of $G$ (resp. $T$). Let $E$ be a $G$-bundle on $X$ with adjoint bundle $E(\mathfrak{g}):=E\times_{G}\mathfrak{g}$. We have a short exact sequence called the Atiyah sequence
[…]
where ${\rm At}(E)$ is the Atiyah algebroid (see [CZ15, A.2] or [Bis10] for instance).
@@
Let $\lambda$ be an element in $k$. A $\lambda$-connection on $E$ is a $\lambda$-twisted splitting of the Atiyah sequence, that is a morphism $\nabla:T_{X}\rightarrow{\rm At}(E)$ such that $q\circ\nabla=\lambda\cdot{\rm id}$. A $\lambda$-connection $\nabla$ is integrable if $[\nabla(u),\nabla(v)]=\lambda\nabla([u,v])$ for any $u,v\in T_{X}$. Such a pair $(E,\nabla)$ is also called a $\lambda$-connection for convenience. The $p$-curvature $\psi_{\nabla}:T_{X}\rightarrow{\rm At}(E)$ of a $\lambda$-connection $\nabla$ is defined by the formula
$\psi_{\nabla}(v)=(\nabla(v))^{p}-\lambda^{p-1}\nabla(v^{[p]})$
for $v\in T_{X}$. If there is no ambiguity, we use the notation $\psi:=\psi_{\nabla}$ for simplicity.
Remark 2.2.
The $p$-curvature $\psi$ of a $\lambda$-connection is $p$-linear and satisfies $q\circ\psi=0$. Then $\psi$ factor through $E(\mathfrak{g})$ and thus can be regarded as a section $E(\mathfrak{g})\otimes F^{*}_{X/k}\Omega_{X^{\prime}}$. A similar discussion was given in [HZ23, §5]. Moreover, a $\lambda$-connection $\nabla$ can be regarded as a map $T_{U}\rightarrow E|_{U}(\mathfrak{g})$ locally on affine open subset $U$.
Remark 2.3.
In this paper, two cases are of most concern: 00-connections ($G$-Higgs bundles) and $1$-connections (flat $G$-bundles).
A 00-connection $\theta$ satisfies $q\circ\theta=0$. Then $\theta$ factors through $E(\mathfrak{g})$. If $\theta$ is integrable, we obtain a Higgs field $\theta:\mathcal{O}_{X}\rightarrow E(\mathfrak{g})\otimes\Omega_{X}$. Therefore, such a pair $(E,\theta)$ is exactly a $G$-Higgs bundle.
A $1$-connection $\nabla$ is exactly a connection. If $\nabla$ is integrable, the pair $(V,\nabla)$ is a flat $G$-bundle. Furthermore, there is an equivalent definition of connections on $G$-bundles, which is also frequently used in this paper. We give it here for the convenience of the reader and refer to [CZ15, A.1] for more details. There is a canonical connection on $\mathcal{O}_{G\times X}$ and denote it by $\nabla_{G}$. A(n) (integrable) $G$-connection $\nabla$ on $E$ is a(n) (integrable) connection $\nabla:\mathcal{O}_{E}\rightarrow\mathcal{O}_{E}\otimes_{\mathcal{O}_{X}}\Omega_{X}$ compatible with the multiplication of $\mathcal{O}_{E}$ such that the following diagram commutes